Hi,
Could you please instruct me how to solve the following problem:
1/2+(2/3 + 3/4) - (4/5 * 5/6) =?

Answer by stanbon(75887)

(1/2)+(2/3 + 3/4) - (4/5 * 5/6) =?
---------------------
Evaluate each of the parentheses:
= (1/2) + [(8+9)/12] - (4/6)
= (1/2) + (17/12) - (2/3)
You need a least common denominator:
lcd = 12
---
Rewrite each fraction with the lcd as its denominator:
= (6/12) + (17/12) - (8/12)
Combine the numerators over the lcd:
= (6 + 17 - 8)/12
Simplify:
= (15/12)
Reduce:
= 5/4
=============
Cheers,
Stan H.